OpenStudy (anonymous):

Find the volume of a rectangular prism if the length is 2x, the width is 4x2, and the height is 2x^2 + x + 7. Use the formula V = l ⋅ w ⋅ h, where l is length, w is width, and h is height, to find the volume.

OpenStudy (anonymous):

You have to multiply V = (2x)(4x²)(2x² + x + 7) Start by multiplying 2x and 4x² Then distribute the result through the parentheses

OpenStudy (anonymous):

but can you multiply those two?... because they have different exponents ?

OpenStudy (anonymous):

you can multiply if they have the same base. (You just can't add or subtract with different exponents.) When multiplying, add the exponents. When subtracting subtract them. For example \[x^5 \times x^6=x^{5+6}=x^{11}\] \[\frac{ x^9 }{ x^4 }=x^{9-4}=x^5\]

OpenStudy (anonymous):

For your problem you have \[2x \times 4x^2=8x^3\]

OpenStudy (anonymous):

okay so it would be (8x^3)(2x^2+x+7) ..?

OpenStudy (anonymous):

right, now distribute

OpenStudy (anonymous):

16x^5+8x^4+56x^3?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

yayy thank you!

OpenStudy (anonymous):

you're welcome